OpenStudy (anonymous):

Simplify (a^8bc^4)^3 · (b^3c)^3

OpenStudy (anonymous):

\[(a^8bc^4)^3 \times (b^3c)^3\]\[(a^{24}b^{3}c^{12}) \times (b^9c^3)\]\[(a^{24}b^{3+9}c^{12+3})\]\[(a^{24}b^{12}c^{15})= Answer\]